OpenStudy (anonymous):

Can anyone teach me how to do these kinds of problems please?: evaluate cos(arctan(3/4))

OpenStudy (therealmeeeee):

ok i will help u

OpenStudy (anonymous):

others: evaluate cos(2pi+arctan4/3)

OpenStudy (therealmeeeee):

sec(t)^2 - tan(t)^2 = 1 sec(t) = 1 / cos(t) sec(arctan(3/4))^2 - tan(arctan(3/4))^2 = 1 sec(arctan(3/4))^2 - (3/4)^2 = (4/4)^2 sec(arctan(3/4))^2 - 9/16 = 16/16 sec(arctan(3/4))^2 = 25/16 sec(arctan(3/4)) = 5/4 1/sec(arctan(3/4)) = 1 / (5/4) cos(arctan(3/4) = 4/5

OpenStudy (therealmeeeee):

do this help u

OpenStudy (therealmeeeee):

@cindylouhoo97 do this help uuu

OpenStudy (campbell_st):

wow I would have drawn a picture and used pythagoras